When my grandmother got married in 1950, both her age and her grandmother's age were respectively equal to the last two digits of their birth years! How old was my grandmother's grandmother when my grandmother got married? WOW!!! Thanks for your help.
Answer by Edwin McCravy(20056)

The grandmother was born in the 1900's and her grandmother would have
been born in the 1800's.

The grandmother was x years old in 1950,
She was born in the year 1900+x
Therefore we have the equation

1950 - her birth year = her age

      1950 - (1900+x) = x
          1950-1900-x = x
                 50-x = x
                   50 = 2x
                   25 = x

So the grandmother was born in 1900+x = 1900+25 = 1925
And she was 25 in 1950.


The grandmother's grandmother was y years old in 1950,
Therefore she was born in the year 1800+y
Therefore we have the equation

1950 - her birth year = her age

      1950 - (1800+y) = y
          1950-1800-y = y
                150-y = y 
                  150 = 2y
                   75 = y

So the grandmother's grandmother was born in 1800+y = 1800+75 = 1875
And she was 75 in 1950.

[Note, we could have just done the second part because the problem
didn't ask anything about the grandmother, but only about the
grandmother's grandmother's age in 1950.]

Edwin